You bought 100 shares of Haslett Helicopters, Inc., in January 2016 for $67.26 per share (assume no commissions, so you spent a total of $6726). The company pays dividends annually, and in 2016 it paid a dividend of 62 cents per share; in 2017 the dividend was 74 cents per share; it 2018 it suspended its dividend indefinitely (i.e., no dividend was paid in 2018 or 2019). You sold the stock in January 2019 at $52.78 (again assume no commissions).

What was your TOTAL return on the stock over the period of January 2016 through January 2019?

What was your ANNUAL return on the stock? Is your calculation based on Geometric or Arithmetic Mean?
Answer #1

Return on stock will be

Dividends and capital gain

Dividends are 0.62 and 0.74

Total dividends is (0.62+0.74)×100 = 136

Capital loss is 100(67.26-52.78) = 948

Return will be

(136-948)/6726 = -12.07%

Annualised return will be

= (1.1207)^1/4 -1 = -2.889%

Our calculation is based in geometric mean